How would I find the equation of an ellipse if I am give the foci(5,0) and (-5,0) and the length of major axis is 12
Answer by lwsshak3(11628) 
 How would I find the equation of an ellipse if I am give the foci(5,0) and (-5,0) and the length of major axis is 12
**
Standard form of equation for ellipse with horizontal major axis:
(x-h)^2/a^2+(y-k)^2/b^2=1,a>b, (h,k) being the (x,y) coordinates of the center.
..
center: (0,0)
length of major axis=12=2a
a=6
a^2=36
..
c=5
c^2=25
..
c^2=a^2-b^2
b^2=a^2-c^2=36-25=11
..
Equation of given ellipse:
x^2/36+y^2/11=1
|
|
|